OpenStudy (dcolley):

Prove the identity: sin^4(x)+cos^4(x)=1-2cos^2(x)+2cos^4(x)

OpenStudy (dcolley):

yes

OpenStudy (experimentx):

\( a^4 + b^4 = (a^2 + b^2)^2 - 2a^2b^2 \) and \( \sin^2x = 1 - \cos^2x\)